OpenStudy (anonymous):

simplify (3x^2 +3x -36) / (x^2-x-6) make sure to list all restrictions

Parth (parthkohli):

lol no one looked at this when I answered xP

OpenStudy (callisto):

(3x^2 +3x -36) / (x^2-x-6) Factor the expression in the numerator first (3x^2 +3x -36) = 3 ( x^2 + x - 12) = 3 ( x^2 +4x -3x - 12) = 3 [ x(x+4) - 3(x+4)] = 3 (x+4)(x-3) Then, factor the denominator. (x^2-x-6) = x^2 - 3x + 2x -6 = x(x-3) + 2(x-3) = (x+2)(x-3) Since it's a fraction, the denominator should NOT be zero. If it is zero, the fraction becomes undefined. So, to find its restriction, put denominator = 0 (x+2)(x-3)=0 x+2 = 0 or x-3 = 0 x = -2 or x=3 Therefore, the restrictions are x ≠-2 AND x≠3 Now, back to the question. \[\frac{(3x^2 +3x -36)}{(x^2-x-6)}\]\[=\frac{ 3 (x+4)(x-3)}{(x+2)(x-3)}\]Cross out the common factor x-3 , you'll get \[=\frac{ 3 (x+4)}{(x+2)}\] Clear?

OpenStudy (anonymous):

sooo youre supposed to put the denominator equal to zero?

OpenStudy (callisto):

Put the denominator equal to zero to get the restriction - Yes

OpenStudy (anonymous):

oh see this makes sense thank you :)

OpenStudy (callisto):

Welcome~~

OpenStudy (anonymous):

so you should always find the restrictions right after you factor and before you completely simplify?

OpenStudy (callisto):

Yes, since you can't cancel 0 as a common factor.

OpenStudy (anonymous):

okay omg thank you so much! Im gonna have another question up soon ;)

OpenStudy (anonymous):

Sometimes it really miff me off that the site doesn't have a edit button :/ \[ \frac {(3x^2 +3x -36)}{ (x^2-x-6)} = \frac{3 (x-3)(x+4)}{(x-3)(x+2)} \] Now about the restriction, cancelling is a confusing word, the better word is divide both of the numerator and denominator by their gcd. In this case it is (x-3), but as we know that division by zero is undefined, hence \((x-3) \neq 0 \implies x \neq 3\). So we get:\[\frac{3 (x+4)}{(x+2)} = \frac{3x+12}{(x+2)}\]\]

OpenStudy (anonymous):

I think it's better to write the final answer as \[ \frac {(3x^2 +3x -36)}{ (x^2-x-6)} =\frac{3x+12}{(x+2)}, \forall x \neq -2 \] after writing all of what I stated before.